OpenStudy (perl):

factor x^4 + 1. over non reals

myininaya (myininaya):

\[(x^2-1)(x^2+1)=(x-1)(x+1)(x-i)(x+i)\]

OpenStudy (perl):

thats wrong

OpenStudy (perl):

x^4 +1 = (x^2)^2 - i ^2 = (x^2 + i)(x^2-i)

myininaya (myininaya):

oops i thought that was a minus sign in between x^4 and 1

myininaya (myininaya):

\[(x^2-i)(x^2+i)\]

OpenStudy (perl):

so im having trouble factoring x^2 + i

myininaya (myininaya):

my bad! its early for me and i don't have my glasses on lol

myininaya (myininaya):

how do you know when to stop factoring?

OpenStudy (anonymous):

\[(x^2 + 1)^2 = (x^2+i)(x^2-i) \]

OpenStudy (perl):

factor it into a+bi , roots

OpenStudy (perl):

youre not there yet , lol

OpenStudy (perl):

this is a delight

OpenStudy (anonymous):

\[x = \pm e^{i \pi/4}\]

OpenStudy (perl):

should be 4 roots

OpenStudy (perl):

,how did you get that evern

OpenStudy (anonymous):

sorry ... x = \pm e^{i (\pi/4 +\pi/2)}

OpenStudy (anonymous):

use euler's form e^ix.. this allows us to calculates roots easily( a root is ^1/k)

OpenStudy (anonymous):

smoething wrong with the renderer ... \[x = \pm e^{i (\pi/4 +\pi/2)}\]

OpenStudy (anonymous):

that was the 2nd set of sols

OpenStudy (anonymous):

that was the 2nd set of sols

OpenStudy (anonymous):

was i any help?

OpenStudy (perl):

yes very much , one sec

OpenStudy (perl):

how did you know it was pi/4

OpenStudy (perl):

we have x = sqrt (i) , -sqrt (i) , sqrt (-i) , -sqrt(-i)

OpenStudy (anonymous):

imagine the complex plane the imaginary axis,which denotes \[ \pi/4 \] is at an angle of 90 degrees to the real axis => 90/2 = 45 =>pi/4 \[i = e^{i \pi/2}\]

OpenStudy (perl):

i dont follow

OpenStudy (perl):

the imaginary axis is pi/2 to the real axis

OpenStudy (anonymous):

yes...and then we sq root it.The square root expressed as an exponential is x^(1/2) hence the entire power bit of the exponent is divided by 2 giving pi/4

OpenStudy (perl):

ok so you did x^2 = e^i(pi/2 +2pik )

OpenStudy (perl):

, for x^2 = i

OpenStudy (anonymous):

yes...and then we sq root it.The square root expressed as an exponential is x^(1/2) hence the entire power bit of the exponent is divided by 2 giving pi/4

OpenStudy (perl):

there are actually an infintie number of solutions, but we are interested in the principal

OpenStudy (anonymous):

yes...i actually skipped out on generalizing 2pi phase factor

OpenStudy (perl):

ok and this follows from the fact that a + b i = r e^i(theta + 2pik) ?

OpenStudy (perl):

, which is euler identity

OpenStudy (perl):

ok so then its just a matter of exponentiation , x = e^[i(pi/2 + 2pik)*1/2]

OpenStudy (anonymous):

yes....again while doing the initial pass over the question (complex analysis) i usually omit the 2pi factor.It necessary for the later steps but it just makes a mess while trying to understand the solution.

OpenStudy (anonymous):

you got it

OpenStudy (perl):

we could have also used the trig form of a + bi = rcos t + i sin t , but this is cleaner

OpenStudy (anonymous):

:)

OpenStudy (perl):

then wed have to use demoivres theorem, exponential is faster

OpenStudy (anonymous):

once again an astute observation with a firm logical reason backing it up

OpenStudy (perl):

yes i dont even remember demoivrs,

OpenStudy (perl):

this is more concise ;)

OpenStudy (perl):

so i have x = +- e^[i(3pi/4)], x = + - e^[i(pi/4)]

OpenStudy (perl):

, why isnt it acceptable to just write the solutions as above

OpenStudy (anonymous):

what grade are you in?

OpenStudy (perl):

x = sqrt (i) , -sqrt (i) , sqrt (-i) , -sqrt(-i)

OpenStudy (perl):

im in college ;)

OpenStudy (perl):

my book gives radical solutions , sqrt2/2 + i sqrt 2 / 2 , etc

OpenStudy (anonymous):

ahh...then there should be no problem in expressing the answer as an exponent

OpenStudy (perl):

no i mean, the book didnt leave it in exponential euler form

OpenStudy (anonymous):

books generally use a computer program to solve their questions and programs rarely try an make stuff easier for people to understand (I should know as i'm doing engineering in computer science)

OpenStudy (anonymous):

books generally use a computer program to solve their questions and programs rarely try an make stuff easier for people to understand (I should know as i'm doing engineering in computer science)

OpenStudy (perl):

so basically they want sqrt i = a + b i , for some a,b , and they got sqrt2/2 + sqrt2/2 i , or something

OpenStudy (perl):

apparently its not good to leave it as i^(1/2) as a solution

OpenStudy (anonymous):

if you insist....here is where we expnad eulers form \[e^{ix} = cos(x) + i sin(x)\]

OpenStudy (perl):

there we go

OpenStudy (perl):

so sometimes we use all three forms

OpenStudy (anonymous):

cos(3pi/4) = -1/sqrt(2) ,sin(3pi/4) = 1/sqrt(2) .....i'm not sure about the exact values just check them on a calculator

OpenStudy (perl):

, err, two forms , etc

OpenStudy (anonymous):

yes sometimes depending on what you need easy additive manipulation or easy multiplicative manipulation you use different forms

OpenStudy (perl):

thx are you good with 3d graphers , i have this website called romanlabs.com , i found a 3d grapher but i cant figure out the code input

OpenStudy (perl):

i also have a calculus problem, Find the derivative of f (x, y) = xy/ (x+ y) at the point (2, -1) in the direction of 6i + 8j . i got it started the partials are

OpenStudy (perl):

< y^2/(x+y)^2, x^2/(x+y^2) >

OpenStudy (perl):

it would be sweet if i could figure out how to use this graphing calculuator

OpenStudy (anonymous):

there are ways of graphing complex numbers (we usually express the imaginary diemensions as colors)

OpenStudy (anonymous):

p.s. if you want to see the function use wolfram alpha (online .....realy good) or gnuplot(offline allows more flexibility but is more complex) ...note both service/programs are free/open source.

OpenStudy (anonymous):

anyway must leave if you have any other related questions feel free to fb me (note: my fb profile has my name misspelled so look for "Erevn")